. The 1977 Super Bowl was played in
The Rose Bowl in Pasadena. In that football game the
Oakland Raiders scored 18 more points than the Minnesota
Vikings. If the total number of points scored was 46, then
what was the final score for the game?

Answer by rapaljer(4671)

Let x = score of Vikings
x+18 = score of Raiders

x + x+18=46
2x+18=46
2x=46-18=28
2x=28
x=14
x+18=32

Check: 14+32=46

The score was 32-14 Raiders

For additional explanation on WORD PROBLEMS, please see my own website. To find my website, click on my tutor name "rapaljer" anywhere in algebra.com. Click on the link there that will take you to the Homepage of my website, which is a subwebsite at Seminole State College of Florida. From my "Homepage", look for the link "Basic, Intermediate, and College Alegebra: One Step at a Time." Click on "Basic Algebra", and look in "Chapter 1" for TWO sections on Word Problems. Of course, you need to see the one called "Factoring Trinomials." Here you will find my own complete non-traditional, user-friendly explanation of the topics that my own students found MUCH easier to understand than the traditional textbooks! Many of my exercises here are explained in great detail in the MATH IN LIVING COLOR pages that go with this curriculum. Best of all, it's all FREE!

I have explanations on nearly every topic in algebra and many from geometry as well, and many videos from my younger days as well. Everything is FREE, and I don't even have advertising on the pages!! If you don't need help with math, then visit my website to see the JOKE PAGE--all clean jokes, safe for the little ears in your family!!!

Dr, Robert J. Rapalje, Retired
Seminole State College of Florida